OpenStudy (anonymous):

Solve the equation: 2x - 20 = 32

OpenStudy (cggurumanjunath):

add 20 on both sides

OpenStudy (jojo4eva):

add 20 to both sides and divide both sides by 2 and you get 26

OpenStudy (cggurumanjunath):

2x - 20 = 32 +20 +20 ========= 2x/2=52/2==>divide by 2 on both sides !

OpenStudy (cggurumanjunath):

x=52/2

OpenStudy (jojo4eva):

so x=26

OpenStudy (jojo4eva):

GOt it @ReaganMHolmes